# Latent Diffusion for Neural Spiking Data

### v1

@@ 4 Related Work
Latent variable models of neural population dynamics   LDNS builds on previous LVMs in neuroscience, which have been extensively applied to infer low-dimensional latent representations of neural spiking data [61, 32, 39, 58, 62, 13, 27, 63, 21, 6] (see [38] for a comprehensive list.) In addition to capturing shared population-level dynamics and dependence on external stimuli [4], LVMs have been extended to allow autoregressive neuron-level (non-Poisson) dynamics [32, 13, 62, 31] or even direct neural interactions [54]. While these methods often have useful inductive biases (e.g., linear dynamical systems [32, 27] or Gaussian process priors [61]), these models are typically not expressive enough to yield realistic neural samples across a range of conditions.
Deep LVMs and other deep learning-based approaches   Variational autoencoders (VAEs) [25] are particularly popular in neuroscience as they allow us to infer low-dimensional dynamics underlying high-dimensional discrete data [63, 16, 46], especially when combined with nonlinear recurrent neural networks [36]. VAEs have been used to infer identifiable low-dimensional latent representations conditioned on behavior [63] and have incorporated smoothness priors using Gaussian Processes to regularize the latent space [16]. However, the generation performance of VAEs is rarely explored in neuroscience. Besides VAEs, generative adversarial networks (GANs [17]) have been proposed to synthesize spiking neural population activity [34, 42]. While GANs produce high-fidelity samples, they are challenging to train reliably and lack a low-dimensional latent space. More recently, transformer-based architectures have also been adapted to model neural activity [5, 60], though often with the focus of accurate decoding of behavior instead of generation of realistic spiking samples, while also lacking an explicit latent space [3]. Lastly, deterministic approaches utilizing RNNs for dynamical systems reconstruction also target low-dimensional latent dynamics underlying neural data [19], but they do not act as probabilistic generative models.
@@ 5 Summary and discussion
First, following other LDMs in the literature, LDNS decouples latent inference and probabilistic modeling of the data, offering flexibility in reusing the trained autoencoder and diffusion model. For the monkey recordings, all diffusion models (unconditional, conditioned on reach angle, and conditioned on hand velocities) operate in the latent space of the same autoencoder, in contrast to existing approaches that require end-to-end retraining for each type of conditioning variable. Second, we show that LDNS autoencoders can be augmented with per-neuron autoregressive dynamics to capture single-neuron temporal dynamics (e.g., refractoriness), which otherwise cannot be captured with population-level shared dynamics. Third, as a result of the length-generalizable autoencoders and diffusion models using S4 layers, LDNS can generate variable-length trials in both the Lorenz example and human cortical recordings—a feature that will be particularly useful in modeling datasets recorded during naturalistic stimuli or behavior.
Altogether, these features enable LDNS to generate realistic neural activity, especially when conditioned on behavioral covariates. In our experiments, we demonstrate that unseen movement trajectories can be used to conditionally generate samples of neural activity, from which we can decode these hypothetical behaviors. This opens up further possibilities for hypothesis generation and testing in silico, potentially enabling stronger links between experimental and computational works.
@@
Broader impact   Realistic spike generation capabilities increase the risk of research manipulation by generating synthetic data that may be difficult to detect. On the other hand, LDNS could be useful for the dissemination of privatized clinical data. Finally, synthetically generated data (conditioned on unseen behavioral conditions) could be useful for augmenting the training of brain-computer interface decoding models.
@@ Acknowledgements
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039), the Human Frontier Science Program (HFSP), and the European Union (ERC, DeepCoMechTome, 101089288). We utilized the Tübingen Machine Learning Cloud, supported by DFG FKZ INST 37/1057-1 FUGG. JK, AS, and JV are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and JV is supported by the AI4Med-BW graduate program. We thank Chethan Pandarinath for providing access to their compute cluster to train AutoLFADS. We thank Christian F. Baumgartner and all Mackelab members for feedback and discussions.
@@
A central component of our autoencoder architecture is the recently introduced structured state space models (S4)[18]. With an input sequence $\mathbf{x}=[x_{1}\ldots x_{T}]\in\mathbb{R}^{T}$ and corresponding output $\mathbf{y}=[y_{1}\ldots y_{T}]\in\mathbb{R}^{T}$, an S4 layer applies the following operation for each timestep
@@
We include temporal information only in the encoder and model the decoder as a lightweight pointwise MLP for the autoencoder (Fig. A1). This allows us to temporally align the latents with the signal, and ensure that no further temporal dynamics are introduced when mapping the latents back into Poisson rates.
@@
We consider four variants of our proposed diffusion model with time-mixing and channel-mixing layers for four different tasks. In all cases, except for the conditioning mechanism, the internal architecture remains the same and is shown in Fig. A1. For diffusion timestep and fixed-length conditioning vector, we shift and scale the inputs and outputs to the time mixing and channel mixing blocks using adaptive instance normalization, as done in Peebles and Xie [37].
@@
We generate samples from LFADS by sampling initial conditions from the Gaussian prior, running the generator RNN forward, and Poisson-sampling spikes from the resulting firing rates.
Appendix A4 Supplementary Figures Monkey Reach task
Figure A2: Correlation matrices for data and samples from Poisson LDNS, concatenated across trials.
Figure A3: Model performance comparison AutoLFADS, LDNS, LDNS with spike history (LDNSsh), and two distinct training data splits as a reference comparison. Mean and standard deviation across 5 folds sampled with replacement.
Figure A4: Population-level and single neuron-level statistics of a) LDNS, b) LDNSsh, c) AutoLFADS, and d) two distinct training data splits.
@@
Figure A7: LDNS captures the correlation structure of the Lorenz dataset a) Both the autoencoder and LDNS-sampled rates capture the ground truth instantaneous correlation structure of the synthetic rates. b) The auto-correlation structure of ground truth and sampled spiking activity matches perfectly in 4 neuron groups, sorted according to correlation strength. Synthetic Lorenz data group $x$ is denoted by $\mathrm{gt}x$, LDNS samples as $\mathrm{ldns}x$. c) The time-lagged cross-correlational structure is also perfectly captured by LDNS in all groups.

### v2

@@ 4 Related Work
Latent variable models of neural population dynamics   LDNS builds on previous LVMs in neuroscience, which have been extensively applied to infer low-dimensional latent representations of neural spiking data [61, 32, 39, 58, 62, 13, 28, 63] (see [38] for a comprehensive list.) In addition to capturing shared population-level dynamics and dependence on external stimuli [4], LVMs have been extended to allow autoregressive neuron-level (non-Poisson) dynamics [32, 13, 62] or even direct neural interactions [54]. While these methods often have useful inductive biases (e.g., linear dynamical systems [32, 28] or Gaussian process priors [61]), these models are typically not expressive enough to yield realistic neural samples across a range of conditions.
Deep LVMs and other deep learning-based approaches   Variational autoencoders (VAEs) [26] are particularly popular in neuroscience as they allow us to infer low-dimensional dynamics underlying high-dimensional discrete data [63, 16, 46], especially when combined with nonlinear recurrent neural networks [36, 21]. VAEs have been used to infer identifiable low-dimensional latent representations conditioned on behavior [63, 21] and have incorporated smoothness priors using Gaussian Processes to regularize the latent space [16]. However, the generation performance of VAEs is rarely explored in neuroscience. Besides VAEs, generative adversarial networks (GANs [17]) have been proposed to synthesize spiking neural population activity [34, 42]. While GANs produce high-fidelity samples, they are challenging to train reliably and lack a low-dimensional latent space. More recently, transformer-based architectures have also been adapted to model neural activity [5, 60], though often with the focus of accurate decoding of behavior instead of generation of realistic spiking samples, while also lacking an explicit latent space [3]. Lastly, deterministic approaches utilizing RNNs for dynamical systems reconstruction also target low-dimensional latent dynamics underlying neural data [19], but they do not act as probabilistic generative models.
@@ 5 Summary and discussion
First, following other LDMs in the literature, LDNS decouples latent inference and probabilistic modeling of the data, offering flexibility in reusing the trained autoencoder and diffusion model. For the monkey recordings, all diffusion models (unconditional, conditioned on reach angle, and conditioned on hand velocities) operate in the latent space of the same autoencoder, in contrast to existing approaches that require end-to-end retraining for each type of conditioning variable. LDNS is also faster to train than AutoLFADS, which requires population-based training to optimize hyperparameters (appendix A3.1). Second, we show that LDNS autoencoders can be augmented with per-neuron autoregressive dynamics to capture single-neuron temporal dynamics (e.g., refractoriness), which otherwise cannot be captured with population-level shared dynamics. Third, as a result of the length-generalizable autoencoders and diffusion models using S4 layers, LDNS can generate variable-length trials in both the Lorenz example and human cortical recordings—a feature that will be particularly useful in modeling datasets recorded during naturalistic stimuli or behavior.
Altogether, these features enable LDNS to generate realistic neural activity, especially when conditioned on behavioral covariates. In our experiments, we demonstrate that unseen movement trajectories can be used to conditionally generate samples of neural activity, from which we can decode these hypothetical behaviors. These generated latent trajectories reflect behavioral information in an interpretable way. Our methodology is general and can be applied to recordings from any brain region, beyond the motor and speech cortex examples shown here. Thus, LDNS opens up further possibilities for hypothesis generation and testing in silico, potentially enabling stronger links between experimental and computational works.
@@
Broader impact   Realistic spike generation capabilities increase the risk of research manipulation by generating synthetic data that may be difficult to detect. On the other hand, LDNS could be useful for the dissemination of privatized clinical data, though we acknowledge the critical importance of protecting data privacy when working with sensitive human participant data. Finally, synthetically generated data (conditioned on unseen behavioral conditions) could be useful for augmenting the training of brain-computer interface decoding models.
@@ Acknowledgements
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039), the Human Frontier Science Program (HFSP), and the European Union (ERC, DeepCoMechTome, 101089288). We utilized the Tübingen Machine Learning Cloud, supported by DFG FKZ INST 37/1057-1 FUGG. JK, AS, and JV are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and JV is supported by the AI4Med-BW graduate program. We thank Chethan Pandarinath for providing access to their compute cluster to train AutoLFADS. We thank Christian F. Baumgartner and all Mackelab members for feedback and discussions. We would like to also thank our reviewers for their insightful comments which improved our paper.
@@
A central component of our autoencoder architecture is the recently introduced structured state space models (S4)[18]. With an input sequence $\mathbf{x}=[x_{1}\ldots x_{T}]\in\mathbb{R}^{T}$ and corresponding output $\mathbf{y}=[y_{1}\ldots y_{T}]\in\mathbb{R}^{T}$, an S4 layer applies the following operation for each timestep –
@@
We include temporal information only in the encoder and model the decoder as a lightweight pointwise MLP for the autoencoder (Supp. Fig. A1). This allows us to temporally align the latents with the signal, and ensure that no further temporal dynamics are introduced when mapping the latents back into Poisson rates.
@@
We consider four variants of our proposed diffusion model with time-mixing and channel-mixing layers for four different tasks. In all cases, except for the conditioning mechanism, the internal architecture remains the same (Supp. Fig. A1). For diffusion timestep and fixed-length conditioning vector, we shift and scale the inputs and outputs to the time mixing and channel mixing blocks using adaptive instance normalization, as done in Peebles and Xie [37].
@@
Appendix A2 Dataset access information
All real-world datasets used in this work are publicly available under open-access licenses. Our work does not involve the collection of new experimental data.
The human BCI dataset is available at https://datadryad.org/stash/downloads/file_stream/2547369 under a CC0 1.0 Universal Public Domain Dedication license. This dataset was originally published in Willett et al. [57]. The data was collected under appropriate ethical oversight, with approval from the Institutional Review Board at Stanford University (protocol #20804).
The monkey reaching dataset (MC_Maze) is available through the DANDI Archive (https://dandiarchive.org/dandiset/000128, ID: 000128) under a CC-BY-4.0 license. This dataset contains sorted unit spiking times and behavioral data from primary motor and dorsal premotor cortex during a delayed reaching task.
@@
We generate samples from LFADS by sampling initial conditions from the Gaussian prior, running the generator RNN forward, and Poisson-sampling spikes from the resulting firing rates. For inclusion of spike history in the observation model of LFADS, we used the same training method and hyperparameter settings as in LDNSsh (appendix A1.3).
Appendix A5 Supplementary baseline comparisons
For an extended baseline comparison, we implemented two additional methods for the task of unconditional generation on the Monkey dataset (Sec. 3.4) – Targeted Neural Dynamical Modeling (TNDM, [21]) and Poisson-identifiable VAE (pi-VAE, [63]). It is important to note that while both TNDM and pi-VAE have demonstrated success in analyzing neural and behavioral data, neither was specifically designed for realistic spike train generation. The architectural choices in our implementation of these methods reflect their original intended applications in neural data analysis rather than generation of neural spiking data. Nevertheless, our comparisons show that LDNS, especially with spike-history, is superior or on par with all other methods (Table 4).
A5.1 Targeted Neural Dynamical Modeling (TNDM)
TNDM [21] is a VAE-based model designed to jointly model neural activity and behavior. TNDM extends LFADS by using an RNN to generate latent dynamics that are mapped to both neural activity and behavioral variables. TNDM separates the latent space into behavior-specific and behavior-independent subspaces to disentangle task-relevant and intrinsic neural dynamics.
For our comparison on the unconditional monkey reach task, we trained TNDM using the architecture and hyperparameters proposed in the original implementation of the paper [^4] Code adapted from Github respository: https://github.com/HennigLab/tndm.. We used 64-dimensional latent dynamics for each of the two sets. These project to a total of 10 latent factors $z$ (5 behavior-specific $z_{r}$ and 5 behavior-independent $z_{i}$), which is the maximum number demonstrated in the original work.
To generate unconditional samples, we sampled initial generator states from a standard normal prior $\mathcal{N}(0,I)$, then generated the latent dynamics and projected into neuron rates the same way as in LFADS. TNDM, performs well in matching real data in spike statistics (Supp. Fig. A13 cyan, Supp. Fig. A16e) and temporal dynamics (Supp. Fig. A15). Overall, we observe that LDNS captures spike statistics better than TNDM, except for the the population spike history count. In all metrics, LDNS augmented with spike history outperforms TNDM on spike statistics.
A5.2 Poisson-identifiable VAE (pi-VAE)
Poisson-identifiable VAE (pi-VAE) [63] is a VAE-based model for count data that ensures identifiability in the latent space. pi-VAE does not model temporal dependencies, instead treating each time point as an independent sample.
We trained pi-VAE on the monkey dataset using the original architecture and hyperparameters [^5] Code adapted from article by Lyndon Duong (2021) – https://www.lyndonduong.com/pivae.. We used a General Incompressible-flow Network as a decoder, with 2 behaviorally relevant dimensions and 2 independent dimensions in the latent space. However, our evaluation context differs significantly from the original paper’s demonstrations: while pi-VAE was initially evaluated on 50ms time bins and straight reaches only, our comparison uses 5ms bins and conditions on angles across all reaches at both middle and end trajectory points. The “label”, or behavior, is presented as a 4-dimensional vector containing the cosine and sine of initial and final reach angles. Since this has a conditional latent space, sampling is performed by sampling angles randomly.
Importantly, sampling from pi-VAE does not introduce any temporal dependence between spike bins within a trial — pi-VAE was not intended to be a generative model of neural spiking data. The lack of temporal modeling in pi-VAE’s is a fundamental limitation for generating realistic spike trains, as evident in our empirical comparisons (Supp. Fig. A13 in yellow, Supp. Fig. A15,A16f). Note that this failure cannot be diagnosed simply from looking at the sampled spiking data (Supp. Fig. A14).
A5.3 Contributions of LDNS in context to LFADS, TNDM and pi-VAE
  • LDNS is designed specifically for the purpose of accurately generating neural spiking data (unconditionally or conditionally)—a task often ignored by other LVMs designed for neural data analysis such as LFADS, pi-VAE, and TNDM.
  • The S4 autoencoder and diffusion model in LDNS are trained in separate stages, offering modularity, while both components naturally account for temporal dependencies (unlike pi-VAE).
  • S4 is autoregressive, similar to other RNN-based models, but empirically we found it to perform better when extending past the training trial length (compared to LFADS, see Sec. A6.1).
  • One feature provided by some neural-behavioral analysis models (such as pi-VAE and TNDM) is an explicit disentangling of neural vs. behavior-relevant latents. While we found LDNS latents contain relevant behavioural information (Fig. 5e-g, Supp. Fig. A17), we did not explicitly supervise the latent space to induce this property.
  • Finally, the spike history-dependent observation model in LDNS is modular and can be optimized post-hoc using rate predictions of any model to improve spike generation quality. We observed this with LDNS as well as LFADS (Table 1).
Table 4: Added Baselines metrics comparison. $D_{KL}$ for the population spike count histogram and RMSE comparisons. Mean and standard deviation across 5 folds sampled with replacement. Bolded entries represent best-performing values for Poisson observation and spike-history observation model, respectively.
Method $D_{KL}$ psch RMSE pairwise corr RMSE mean isi RMSE std isi
[…]
@@
Figure A4: LDNS captures the correlation structure of the Lorenz dataset a) Both the autoencoder and LDNS-sampled rates capture the ground truth instantaneous correlation structure of the synthetic rates. b) The auto-correlation structure of ground truth and sampled spiking activity matches perfectly in 4 neuron groups, sorted according to correlation strength. Synthetic Lorenz data group $x$ is denoted by gt $x$, LDNS samples as ldns $x$. c) The time-lagged cross-correlational structure is also perfectly captured by LDNS in all groups.
A6.1 Length generalization of LFADS on Lorenz
To analyze whether LFADS [36] exhibits similar length generalization properties as LDNS, we trained an AutoLFADS model on the Lorenz dataset (256 bins). We used the same architecture as the Monkey dataset, with 40-dimensional latent dynamics. We sampled initial conditions from the LFADS prior, then generated dynamics for both the original 256 steps and for an extended duration of 16×16$\times$16 × the training length.
Qualitatively, we observed that while LFADS produced trajectories resembling the attractor dynamics of the ground truth Lorenz system (Supp. Fig. A5a, across various dimension combinations), these trajectories often diverged when run for longer intervals (Supp. Fig. A5b). However, the system eventually returned to typical dynamics.
Furthermore, when generating for extended durations, we observed that the mean population firing rates sometimes reached extreme values in some samples (Supp. Fig. A6), though they eventually returned to typical ranges. This behavior was not observed in LDNS samples, suggesting that bidirectional generation in the diffusion model provides more stability in variable-length generation.
Figure A5: Length generalization of LFADS on Lorenz Different projections of a 40-dimensional latent space from LFADS trained on the Lorenz system. Trajectories are compared between a) the original length and b) 16 times the original length using sampled initial conditions. For comparison with LDNS length generalization, see Fig. 2c
Figure A6: Mean population firing rates for eight different samples, shown for both the original length (pink) and 16×16$\times$16 × the original length (purple).
Appendix A7 Supplementary Figures Human BCI
Figure A7: Visual comparison of different sampled spiking data from LDNS, with five samples from the true dataset.
Figure A8: First two principal components (PCs) of smoothed spikes from true data (left) and model samples (right). Each line represents one sampled trial. Spikes were smoothed using a Gaussian filter with a window of 160ms prior to extracting the PCs.
Figure A9: Correlation matrices for real spiking data from human and LDNS, comparing the autoencoder-inferred (ae) correlation (sampled from reconstructed rates) and correlation of sampled spikes (diff). The deviations from the data (gt) already arise at the autoencoder stage.
Appendix A8 Supplementary Figures Monkey Reach task
Figure A10: Correlation matrices for real spiking data and samples from Poisson LDNS, concatenated across trials, comparing the autoencoder-inferred (ae) correlation (sampled from reconstructed rates) and correlation of sampled spikes (diffusion + ae). Most deviations from the data (gt) already arise at the autoencoder stage.
Figure A11: Visual comparison of different sampled spiking data from LDNS with five samples from the real dataset.
Figure A12: Equipping LFADS with spike history Auto-correlation of data, LFADS samples with Poisson observations (left) and LFADSsh samples (with spike history), grouped according to correlation strength.
Figure A13: Performance comparison with additional baselines pi-VAE [63], TNDM [21], LFADS [36], LFADS with spike history (LFADSsh), LDNS and LDNS with spike history (LDNSsh). Mean and standard deviation across 5 folds sampled with replacement.
Figure A14: Visual comparison of sampled spiking data from LDNS and all baselines with real data.
Figure A15: Comparing principal components of smoothed sampled spikes a) First two principal components (PCs) of smoothed spikes from model samples. Each line represents one sampled trial. Spikes were smoothed using a Gaussian filter with a window of 40ms prior to extracting the PCs. The PCs were fit using real data. Since pi-VAE does not account for temporal dynamics, it does not show any temporal structure in the PCs. b) Power spectral density (PSD) of PC1, plotted for model vs. data (in grey). c) Mean squared error of median PSD between model samples and data for PC1 (left) and PC2 (right). LDNS and LDNSsh perform the best here, with pi-VAE showing large errors.
Figure A16: Population-level and single neuron-level statistics of a) LDNS, b) LDNSsh (with spike history), c) LFADS, d) LFADSsh (with spike history), e) TNDM, and f) pi-VAE.
Figure A17: Comparison of LDNS latent space trajectories of inferred and conditionally sampled latents a) Straight reaches from the Monkey reach test set. b) LDNS sampled latents (velocity-conditioned on reaches shown in a). c) PCs of LDNS sampled latents. d) Autoencoder-inferred latents of corresponding neural activity for the reaches shown in a). e) PCs of autoencoder-inferred latents.